OpenStudy (anonymous):

Evaluate the square root of one hundred plus two squared minus seven minus the square root of thirty-six.

OpenStudy (anonymous):

HELP PLEASE?!

OpenStudy (anonymous):

I'm pretty sure it's sqrt(97)-6 ..

OpenStudy (anonymous):

\[\sqrt{100}-2^2-7-\sqrt{36}\]

OpenStudy (goldphenoix):

\[\large \large \sqrt{100+2^2-7} - \sqrt{36}\]

OpenStudy (goldphenoix):

What is the square root of 36?

OpenStudy (anonymous):

6

OpenStudy (goldphenoix):

Good job. \[\large \large \sqrt{100+2^2-7} - 6\] Now the number inside the square root, you simplify it. So you have 100+ 2^2 - 7. Now simplify that. What do you get?

OpenStudy (anonymous):

i'm not too sure..

OpenStudy (goldphenoix):

Hmm. Wait I think I mest up.

OpenStudy (anonymous):

ook

OpenStudy (goldphenoix):

Hmm. I believe it's suppose to be the way doulike wrote. Except with one minor changes. \[\large \large \sqrt{100} + 2^2 - 7 - \sqrt{36}\] We know that square root of 36 is 6. Now what's the square root of 100?

OpenStudy (anonymous):

10

OpenStudy (goldphenoix):

Great. Now it should look like: \[\large \large 10 + 2^2 - 7 - 6\] 2^2 give us?

OpenStudy (anonymous):

i don't get it :/

OpenStudy (goldphenoix):

\[\large \large 2^2 = ?\]

OpenStudy (anonymous):

4

OpenStudy (goldphenoix):

Great. Now 10 + 4 - 7 - 6 = ?

OpenStudy (anonymous):

1

OpenStudy (goldphenoix):

Great. So the answer is 1. I hope this help! :)

OpenStudy (anonymous):

so it's 1? thank you so much!

OpenStudy (goldphenoix):

No problem. :)